# SAI3D: Segment Any Instance in 3D Scenes

### 17 Dec 2023 v1

@@ 4.3 Ablation and Analysis
Table 4: The ablation studies of the effectiveness of our designs. The experiments are conducted on ScanNet200 dataset.
Effect of our designs on the region growing algorithm.
We analyze the three key design choices of our region growing algorithm: the adoption of 3D superpoint primitive, Multi-level merging criteria, and Progressive growing. The results of ablation studies are shown in Table 4. We find that omitting the 3D superpoints and using a classic point-based region-growing approach leads to a noticeable drop in performance. The incorporation of multi-level merging criteria and the use of progressive growing both contribute to enhancing the algorithm’s robustness, particularly in dealing with the noise present in 2D segmentation masks, yielding better results respectively. When all the proposed designs are combined, there is an approximately 50% improvement in performance compared to the original, basic version of the algorithm, underlining the effectiveness of our design choices of the algorithm.
Figure 8: The ablation studies of the correlation between the performance and the number of 2D images. The experiments are conducted on ScanNet200 dataset.
Effect of the number of 2D images.
As both our approach and the concurrent method SAM3D leverages 2D image masks, we thus study the method’s robustness against the number of images. Among all the images of a 3D scene, we vary the proportion of views from 1% to 20% and report the performance curve in Fig. 8. Our approach progressively improves the 3D segmentation performance when more views are provided, while SAM3D fails to aggregate useful information when the proportion of views becomes larger than 5%. More importantly, our approach achieves an AP score larger than 20 with only 1% of views, outperforming the best performance achieved by SAM3D. This highlights the effectiveness of our 3D primitive-based region-growing algorithm.
@@ 5 Conclusion
We have introduced SAI3D, a novel approach for zero-shot 3D instance segmentation. Our method offers an efficient alternative to supervised methods which are confined to specific datasets and hence inhibit broader applicability in open-world scenarios. The technique we presented builds upon the power of 2D and 3D segmentation techniques, and a novel progressive region-growing algorithm, which smartly merges the 3D superpoints into object masks that define the final 3D instance segmentation.
Through our evaluation on ScanNet and ScanNet++, we demonstrated that our approach significantly outperforms prior unsupervised methods like SAM3D in class-agnostic segmentation, and even surpass fully-supervised Mask3D models trained with 3D annotations on the more challenging ScanNet++ dataset. This success underscores the potential of leveraging geometric primitives and multi-view consistency to achieve high-quality instance segmentation in complex 3D scenes, without using 3D labeled data.
Limitations.
Our method fundamentally relies on accurate 2D segmentation results. This constitutes the main limitation of our method. As our approach builds a scene graph based on 3D primitives and computes an affinity matrix based on 2D masks, incorrect 2D masks naturally result in unreliable affinities, thus affecting the fine segmentation results. Though we design a simple voting scheme to avoid this error, designing a more advanced 2D mask aggregation mechanism should be a promising direction.
Another limitation of our approach is the running speed. As our approach involves aggregating 2D mask segmentation from images, the total processing time could linearly scale with the number of images. This makes it difficult to apply our method in large-scale scenes. Designing a more efficient algorithm without iterating over all images is another important future direction.
@@ Appendix A Implementation Details
Following the baselines [38, 40], we evaluate the numerical results on the validation set of the datasets. We follow UnScene3D [38] to implement class-agnostic instance segmentation, where all object categories are treated equally and only the mask AP values are considered. For 2D foundation models, we choose SAM-HQ [17] on ScanNet++ dataset, and Semantic-SAM [23] on ScanNet dataset for better granularity control.
Code release.
We will release the code upon acceptance.
@@ Appendix B Additional Qualitative Results
Figure 10: Visual results of 3D instance segmentation on ScanNet dataset.

### 24 Mar 2024 v2

@@ 4.3 Ablation and Analysis
Table 5: The ablation studies of the effectiveness of our designs. The experiments are conducted on ScanNetV2 dataset.
@@
Effect of our designs on the region growing algorithm. We analyze the three key design choices of our region growing algorithm: the adoption of 3D superpoint primitive, Multi-level merging criteria, and Progressive growing. The results of ablation studies are shown in Table 5. We find that omitting the 3D superpoints and using a classic point-based region-growing approach leads to a noticeable drop in performance. The incorporation of multi-level merging criteria and the use of progressive growing both contribute to enhancing the algorithm’s robustness, particularly in dealing with the noise present in 2D segmentation masks, yielding better results respectively. When all the proposed designs are combined, there is an approximately 50% improvement in performance compared to the original, basic version of the algorithm, underlining the effectiveness of our design choices of the algorithm.
Figure 8: Ablation studies on the correlation between performance and the number of 2D images. We define coverage and cross coverage as the percentage of points observed at least once and more than 10 times, respectively. The experiments are conducted on ScanNetV2 dataset.
Effect of the number of 2D images. As both our approach and the concurrent method SAM3D leverages 2D image masks, we thus study the method’s robustness against the number of images. Among all the images of a 3D scene, we vary the proportion of views from 1% to 20% and report the performance curve in Fig. 8. Our approach progressively improves the 3D segmentation performance when more views are provided, while SAM3D fails to aggregate useful information when the proportion of views becomes larger than 5%. More importantly, our approach achieves an AP score larger than 20 with only 1% of views, outperforming the best performance achieved by SAM3D. This highlights the effectiveness of our 3D primitive-based region-growing algorithm.
@@ 5 Conclusion
We introduce SAI3D, a novel approach for zero-shot 3D instance segmentation. Our method offers an efficient alternative to supervised methods which are confined to specific datasets and hence inhibit broader applicability in open-world scenarios. The technique we presented builds upon the power of 2D and 3D segmentation techniques, and a novel progressive region-growing algorithm, which smartly merges the 3D superpoints into object masks that define the final 3D instance segmentation.
Through our evaluation on ScanNet, ScanNet++ and Matterport3D datasets, we demonstrate that our approach significantly outperforms prior unsupervised methods like SAM3D in class-agnostic segmentation, and even surpass fully supervised Mask3D models trained with 3D annotations on the more challenging ScanNet++ dataset. This success underscores the potential of leveraging geometric primitives and multi-view consistency to achieve high-quality instance segmentation in complex 3D scenes, without using 3D labeled data.
Limitations. Our method fundamentally relies on accurate 2D segmentation results and reliable 2D-3D alignment. As our approach builds a scene graph based on 3D primitives and computes an affinity matrix based on 2D masks through 2D-3D lifting, incorrect 2D masks or camera poses naturally result in unreliable affinities, thus affecting the fine segmentation results. Though several techniques are designed to better leverage geometry priors and multi-view consensus, designing a more advanced 2D mask aggregation mechanism should be a promising direction.
Another limitation is the running speed. As our approach involves aggregating 2D mask segmentation from images, the total processing time linearly scale with the number of images making it difficult to apply in large-scale scenes. Designing a more efficient algorithm without iterating over all images is another important future direction.
@@ Appendix A Implementation Details
Datasets.
To balance the performance and the efficiency, we sample proportions of images for different datasets. We use $5\%$ views for ScanNet++ dataset, $20\%$ views for both ScanNetV2 and ScanNet200 dataset, and all images in Matterport3D since it is a sparsely-scanned dataset.
Following the common practice [43, 45, 29], we ignore the instances with semantics of “wall” and “floor” for ScanNetV2 and ScanNet200 datasets. For Matterport3D dataset, we consider the most frequent 160 classes provided in [34] and ignore “wall”, “floor” and “ceiling”. For ScanNet++ dataset, we use all GT labels provided for the instance segmentation task (in “instance_classes.txt”). IPhone images are served as our input to better reflect everyday cases.
@@
Following the baselines [41, 43, 45, 34], we evaluate the numerical results on the validation set for ScanNetV2, ScanNet200 and ScanNet++ datasets and on the test set for Matterport3D dataset. We follow UnScene3D [41] to implement class-agnostic instance segmentation, where all object categories are treated equally and only the mask AP values are considered. We set all the confidence scores as 1.0, the same as [45]. For 2D foundation models, we choose SAM-HQ [18] for ScanNet++ dataset, and Semantic-SAM [24] for ScanNetV2, ScanNet200 and Matterport3D datasets for better granularity control.
@@ Appendix B Additional Qualitative Results
Figure 10: Visual results of 3D instance segmentation on ScanNetV2 dataset.